Answer to #1. 2& 3 please. Thanks.


Answer to #1.2& 3 please. Thanks. Answer to #1.


Answer to #1.2& 3 please. Thanks.2& 3 please. Thanks.
Answer #1

MOLAR ENTHALPY OF WAX:

The formula for the candle wax is C25H52

and the accepted value for the heat of combustion is 15MJ / mol.

  Assumption is heat given off by candle = heat captured in the water.

heat captured in water = (4.184 J/g-degC)(1750 g)(24-21)degC
= 22000 Joules

Now find moles of candle burned.
mass burned = 5.88-5.24 = 0.64 g
molar mass of C25H52 = 353 g/mol
moles wax burned = 1.8 x 10^-3 mol

Then convert to heat per mole:


Heat per mole = 22000 J/1.8 x 10^-3 mol = 12,000,000 J/mol

OR

,( 12 MJ/mol)

Initial mass of candle: 5.88g

Final mass of the candle=5.24g

initial temperature of the water=21

final temperature of the water=24

mass of the water =1.75L